OpenStudy (anonymous):

Four pine seedling and 3 fir seedlings cost 29.00. Two pine seedlings and 7 fir seedlings cost $42.00 Determine the cost of each type of seedling.

OpenStudy (anonymous):

p=price of pine f=price of fir So: 4p+3f=29 2p+7f=42

OpenStudy (anonymous):

Now its a system of equations you can solve

OpenStudy (anonymous):

Yes, and what do i do from there?

OpenStudy (anonymous):

Well to solve it you wanna eliminate one of the variables when you add the equations together, so lets multiply the bottom one by -2 4p+3f=29 -4p-14f=-84 Add the two equations together: -11f=-55 Now you can solve for f

OpenStudy (anonymous):

Why are we mulitpying the bottom by two?

OpenStudy (anonymous):

It's -2 and thats so that when we add the two equations together, one of the variables will cancel out, that way well have only one variable so we can solve

OpenStudy (anonymous):

Ohhh, I got it! I was so lost! Thanks! Can I do one more problem with you and can you see if I did it correctly?

OpenStudy (anonymous):

Don't you wanna finish this one? What'd you get for the price of a fir seedling?

OpenStudy (anonymous):

Sure, okay, give me a couple seconds and I'll send it your way. :)

OpenStudy (anonymous):

Okay, I got a weird answer...I got 11 for the pine seedlings and -5 for fir seedlings. What did I do wrong?

OpenStudy (anonymous):

The equation for the fir was -11f=-55 divide both sides by -11 and you get 5. So that was right except for the negative sign. Since you know f is 5, you plug that into one of the original equations and solve for p

OpenStudy (anonymous):

Okay, I got 3,50 for the only one.Correct? :)

OpenStudy (anonymous):

Yep so fir=$5 and pine=$3.50

OpenStudy (anonymous):

@CalebBeavers

OpenStudy (anonymous):

Got it?

OpenStudy (anonymous):

Yessir! Thanks! :)

OpenStudy (anonymous):

You're welcome